# TFPI: Thinking-Free Policy Initialization Makes Distilled Reasoning Models More Effective and Efficient Reasoners

### v1

@@ 1 Introduction
Reasoning is a fundamental aspect of human cognition, and equipping artificial intelligence (AI) with strong reasoning capabilities is critical for its deployment and applications (Morris et al., 2023; Huang et al., 2024; Xu et al., 2025c). Progress in pretraining (Shao et al., 2024; Yang et al., 2024; Chen et al., 2025b), supervised fine-tuning (SFT) (Yu et al., 2023; Xu et al., 2024; Tong et al., 2024; Ye et al., 2025; Muennighoff et al., 2025), rigorous evaluation (Rein et al., 2024; Phan et al., 2025; Xu et al., 2025b), reinforcement learning (RL) (Jaech et al., 2024; Guo et al., 2025) has significantly enhanced the reasoning abilities of large language models (LLMs). Among these, RL with verifiable rewards (RLVR) stands out as an effective approach that enables large language models (LLMs) to generate long Chains-of-Thought (CoT) (Wei et al., 2022) spontaneously, and empowers them with unprecedented performance on challenging reasoning tasks. Thus, these RLVR-trained LLMs are termed as long-CoT LLMs, slow-thinking LLMs, or large reasoning models (LRMs).
Compared with initializing from a base LLM, starting from an SFT-distilled LRM typically yields better results and accelerates convergence in RLVR (Guo et al., 2025; An et al., 2025). However, SFT-distilled LRMs often produce excessively long responses during the rollout stage of RLVR, which necessitates a large training context length for RLVR. Using such large training contexts also incurs substantial computational costs. A common mitigation strategy is multistage RLVR, which begins with a relatively “short” context and gradually increases the training length (Luo et al., 2025b; An et al., 2025). Nonetheless, Zeng et al. (2025a) argue that multistage training might cause irreversible performance degradation. Moreover, even multistage training demands significant computational resources. For instance, training a 4B model while progressively increasing the maximum context length from 40K to 48K to 52K tokens requires approximately 8K H800 GPU hours (An et al., 2025). These limitations underscore the need for more efficient RLVR training methods.
@@ 2 Preliminary
Proximal Policy Optimization (PPO). PPO (Schulman et al., 2017) constrains the policy update within a proximal region of the old policy $\pi_{\theta_{\text{old}}}$ through the clipping mechanism. Specifically, PPO employs the following objective (we omit the KL regularization term hereinafter for brevity):
@@
Group Relative Policy Optimization (GRPO). GRPO (Shao et al., 2024) bypasses the need for the value model by computing the relative advantage of each response within a group of responses to the same query. Specifically, GRPO optimizes $\mathcal{J}_{\text{GRPO}}(\theta)=\mathbb{E}_{x\sim\mathcal{D}}\left[\mathcal{J}_{\text{GRPO}}(\theta,x)\right],$ where:
@@
Numerous variants have been proposed to improve GRPO, and any RLVR algorithm can be applied to our proposed TFPI stage. In all our experiments, we adopt one widely used variant, DAPO (Yu et al., 2025), as our RLVR algorithm for fair comparison. Details are provided in Appendix A.1.
@@ 3.1 Thinking-Free Mode Enables More Efficient Reasoning
To generate a response $y\sim\pi_{\theta}(\cdot\mid x)$ for a query $x$ using an SFT-distilled LRM $\pi_{\theta}$, the query is typically formatted with a chat template. Template 1 illustrates the template adopted by the Qwen model family (Yang et al., 2025a). We define ThinkingFree as an operator that transforms an input query $x$ into a modified query $x^{\prime}=\textit{ThinkingFree}{}(x)$, in which the thinking content is explicitly omitted. Under this transformation, response generation follows $y\sim\pi_{\theta}(\cdot\mid x^{\prime})$. This mechanism provides explicit control over whether reasoning content is present or absent in the generated output (see Template 2). Additional examples using other chat templates are provided in Appendix A.2. Hereinafter, we use $x$ to denote a query formatted with the thinking template (e.g., Template 1), and $x^{\prime}$ or $\textit{ThinkingFree}(x)$ to denote the corresponding thinking-free version (e.g., Template 2).
@@
Figure 2: Results of the meta-experiment on the ThinkingFree operation. Left: Average output tokens in thinking mode and ThinkingFree mode on AIME25. Right: Evolution of avg@32 and average output tokens on AIME24 with thinking-mode evaluation over training steps.
@@ 3.2 Thinking-Free Training Is Beneficial to Slow-Thinking
To train an SFT-distilled LRM, the training response length should not be too short (Setlur et al., 2025), as this is detrimental to testing performance (An et al., 2025). As evidenced by the dotted line in Figure 2 (Right), training Qwen-3-4B with a 4K response length using DAPO indeed leads to a substantial drop in performance on AIME25. Given that using the ThinkingFree variant for inference can significantly reduce token consumption, we pose an audacious question: can we apply the ThinkingFree operation to all input queries during the rollout stage of RLVR? Moreover, will this approach be beneficial to preserving the original slow-thinking capability of the trained LLM?
@@ 3.3 Thinking Free Policy Initialization
Hu et al. (2025) introduce a pre-pretraining stage using formal language to accelerate convergence of pretraining, while Wang et al. (2025b) propose a mid-training stage between pretraining and RL-zero to facilitate RLVR. Inspired by these works, we ask: can a dedicated stage for SFT-distilled LRMs improve the efficiency and effectiveness of standard RLVR scaling? ThinkingFree RL, which enhances slow-thinking within short training context windows, requires substantially less compute than standard RLVR. Initializing RLVR with a ThinkingFree RL policy could therefore reduce rollout tokens while achieving stronger downstream performance. We term this step as Thinking Free Policy Initialization (TFPI), a stage preceding standard RLVR for SFT-distilled LRMs that aims to lower rollout costs, raise the ceiling of reasoning ability, and accelerate convergence.
@@
where $r(x^{\prime},y)=r(x,y)$ since the ThinkingFree operator does not alter the ground-truth answer of the original problem. In our experiments, we instantiate RLVR with DAPO, i.e., $\mathcal{J}_{\text{RLVR}}(\theta)=\mathcal{J}_{\text{DAPO}}(\theta)$ (see eq. 5). The details of DAPO (Yu et al., 2025) algorithm is given in Appendix A.1
@@ 4.1 Experimental Settings
Baselines. We compare TFPI with direct RLVR training from an SFT-distilled LRM (“Direct RL”) under matched total training compute (Table 1). To assess TFPI as a pre-RLVR stage, we run “TFPI + RL” with similar compute and compare against “Direct RL” (Table 2), also including competitive LRMs of the same size, such as Polaris (An et al., 2025), DeepScaleR (Luo et al., 2025b), Skywork-OR1 (He et al., 2025), and so on. For efficiency analysis, we compare with several RL-based efficient reasoning baselines (Table 3) under the same evaluation for fair comparison, including TLMRE (Arora & Zanette, 2025), AdaptThink (Zhang et al., 2025b), AutoThink (Tu et al., 2025), Laser (Liu et al., 2025a), L1Max (Aggarwal & Welleck, 2025), and ThinkLess (Fang et al., 2025).
Evaluation Details. Our evaluation benchmarks cover: ❶ Math Reasoning: AIME24/25 and BeyondAIME (ByteDance-Seed, 2025). ❷ Multi-Task Reasoning: GPQA-Diamond (Rein et al., 2024). ❸ Code Generation: LiveCodeBench (Jain et al., 2024). ❹ Instruction Following: IFEval (Zhou et al., 2023). ❷ ❸ ❹ are out-of-domain evaluation. Following Guo et al. (2025), we generate multiple outputs (ranging from 4 to 32 depending on the size of the test set) per problem and report pass@1 accuracy. Note that ❶ is in-domain evaluation, and For IFEval, we report the strict prompt accuracy. All evaluation scripts are adapted from the DeepscaleR codebase (Luo et al., 2025b), with detailed decoding parameters provided in Appendix C.3.
@@ 5.1 Why TFPI Leads to Better Thinking-Mode Inference?
Behavioural level: The learned verification behaviours after </think> during TFPI can generalize to the slow-thinking verification occurring within the <think> and </think>. The blue lines in Figure 3 show the ratio of verification steps (i.e., the number of verification steps divided by the total number of steps; see Appendix D for details) for the training set (thinking-free mode) and AIME 25 dataset (thinking mode). We observe that the verification ratio exhibits a similar trend: a rapid drop in Stage 1, followed by steady growth in Stage 2, and a sharp increase in Stage 3. Notably, the sharp decline in Stage 1 resembles an information compression process. In Stages 2 and 3, the model begins to explore more extensively (Figure 3 Right), which may explain why TFPI achieves superior performance. As verification is believed to be vital for slow-thinking reasoning (Setlur et al., 2025), the observed generalization of verification behavior suggests a transfer from thinking-free training to inference in thinking mode.
@@ 6 Related Work
LRMs & Efficiency RLVR. RLVR has enabled the development of numerous high-performing large reasoning models (LRMs) (Yang et al., 2025a; Zeng et al., 2025a; An et al., 2025), inspiring research on model behaviors (Liu et al., 2025b; Wang et al., 2025a), novel algorithms (Liu et al., 2025b; Zheng et al., 2025), multimodal extensions (Meng et al., 2025; Xiao et al., 2025; Wang et al., 2025c), tool integration (Jin et al., 2025; Feng et al., 2025; Li et al., 2025a; Xue et al., 2025; Team et al., 2025a), and other related directions (Xu et al., 2025a; Zhang et al., 2025a; Chen et al., 2025c; Zhang et al., 2025c). RLVR can be applied directly to base LLMs (“RL zero”) (Zeng et al., 2025b; Guo et al., 2025) or initialized from SFT-distilled long-CoT models, the latter typically yielding stronger results (Luo et al., 2025b; An et al., 2025). A major challenge for RLVR is the cost of training with long contexts, as longer outputs are often necessary for harder tasks (Shrivastava et al., 2025; Zeng et al., 2025a), consuming high computational resources. Multi-stage RLVR mitigates this by starting with shorter contexts and gradually extending them (Luo et al., 2025b; An et al., 2025; He et al., 2025), while algorithmic approaches modify GRPO to reduce length bias (Yu et al., 2025; Liu et al., 2025b; Wang et al., 2025a). Orthogonal to these strategies, we introduce TFPI as a lightweight stage before RLVR, improving efficiency and strengthening the slow-thinking mode at inference with minimal training cost, thus facilitating more effective subsequent RLVR.
Efficient Reasoning. To address the issue of overthinking (Chen et al., 2024; Sui et al., 2025), considerable efforts have been made, including prompt-based methods (Muennighoff et al., 2025; Yang et al., 2025c; Fu et al., 2025a; Chen et al., 2025a; Fu et al., 2025b), SFT-based approaches (Kang et al., 2025; Ma et al., 2025; Munkhbat et al., 2025; Luo et al., 2025a), and RL-related designs. RL-related approaches can be further categorized into length-based reward shaping (Team et al., 2025b; Aggarwal & Welleck, 2025; Arora & Zanette, 2025; Liu et al., 2025a), integration of fast and slow thinking (Fang et al., 2025; Zhang et al., 2025b; Lou et al., 2025; Tu et al., 2025; Jiang et al., 2025; Zhang et al., 2025d), and thinking budget control (Li et al., 2025b; Hammoud et al., 2025; Wen et al., 2025). These methods primarily trade accuracy for efficiency and rely on specialized reward functions or training strategies to encourage more efficient reasoning. In contrast, our proposed TFPI naturally yields even more efficient LRMs without specialized rewards or training designs.
@@ A.1 RLVR Algorithms
Numerous variants have been proposed to improve GRPO. For example, DAPO (Yu et al., 2025) introduces token-level normalization and dynamic sampling; Dr GRPO (Liu et al., 2025b) removes length bias to prevent incorrect responses from growing longer over time; and Wang et al. (2025a) train selectively on forking tokens (see also Section 6). Our TFPI is orthogonal to these RLVR algorithms. That is to say, any RLVR algorithm can be applied to our proposed TFPI stage. To mitigate the effect of RLVR algorithms, we employ DAPO as our RLVR algorithm in all experiments for fair comparison.
@@
Building on GRPO, DAPO (Yu et al., 2025) introduces a clip-higher mechanism, incorporates dynamic sampling and applies a token-level policy gradient loss. The objective is given by $\mathcal{J}_{\text{DAPO}}(\theta)=\mathbb{E}_{x\sim\mathcal{D}}\left[\mathcal{J}_{\text{DAPO}}(\theta,x)\right],$ where:
@@ A.2 ThinkingFree Operation
In Section 3.1, we have shown chat templates for both the original query $x$ and its thinking-free version $x^{\prime}=\textit{ThinkingFree}(x)$ for Qwen models. Here, we showcase one additional example under the DeepSeek (Guo et al., 2025) template as below.
@@ B.1 Token Consumption of ThinkingFree
For DS-1.5B, we use the decoding parameters suggested by Guo et al. (2025) in thinking mode: temperature = 0.6, top-$p$ = 0.95, and top-$k$ = 1-1. For Qwen3-4B, we adopt the recommended settings from Yang et al. (2025a) in thinking mode: temperature = 0.6, top-$p$ = 0.95, and top-$k$ = $20$. For ThinkingFree, we set temperature = 0.7, top-$p$ = 0.8, and use both top-$k$ = 1-1 and top-$k$ = $20$. The maximum output length is fixed at 32K tokens for both modes. For evaluation, we sample 32 generations per query on AIME 2025 and report the average number of output tokens. The parameters are given in Table 4 and the results are shown in Figure 2 (Left).

### v2

@@ 1 Introduction
Reasoning is a fundamental aspect of human cognition, and equipping artificial intelligence (AI) with strong reasoning capabilities is critical for its deployment and applications (morris2023levels; huang2024olympicarena; xu2025ugmathbench). Progress in pretraining (shao2024grpo; yang2024qwen25math; chen2025gpas), supervised fine-tuning (SFT) (metamath2023yu; xu2024egsm; dartmath2024tong; ye2025limo; muennighoff2025s1), rigorous evaluation (rein2024gpqa; phan2025humanity; xu2025ugphysics), reinforcement learning (RL) (jaech2024openaio1; guo2025deepseekr1) has significantly enhanced the reasoning abilities of large language models (LLMs). Among these, RL with verifiable rewards (RLVR) stands out as an effective approach that enables large language models (LLMs) to generate long Chains-of-Thought (CoT) (CoT2022Wei) spontaneously, and empowers them with unprecedented performance on challenging reasoning tasks. Thus, these RLVR-trained LLMs are termed as long-CoT LLMs, slow-thinking LLMs, or large reasoning models (LRMs).
Compared with initializing from a base LLM, starting from an SFT-distilled LRM typically yields better results and accelerates convergence in RLVR (guo2025deepseekr1; Polaris2025). However, SFT-distilled LRMs often produce excessively long responses during the rollout stage of RLVR, which necessitates a large training context length for RLVR. Using such large training contexts also incurs substantial computational costs. A common mitigation strategy is multistage RLVR, which begins with a relatively “short” context and gradually increases the training length (deepscaler2025; Polaris2025). Nonetheless, zeng2025glm4_5 argue that multistage training might cause irreversible performance degradation. Moreover, even multistage training demands significant computational resources. For instance, training a 4B model while progressively increasing the maximum context length from 40K to 48K to 52K tokens requires approximately 8K H800 GPU hours (Polaris2025). These limitations underscore the need for more efficient RLVR training methods.
@@ 2 Preliminary
Proximal Policy Optimization (PPO). PPO (schulman2017ppo) constrains the policy update within a proximal region of the old policy $\pi_{\theta_{\text{old}}}$ through the clipping mechanism. Specifically, PPO employs the following objective (we omit the KL regularization term hereinafter for brevity):
@@
Group Relative Policy Optimization (GRPO). GRPO (shao2024grpo) bypasses the need for the value model by computing the relative advantage of each response within a group of responses to the same query. Specifically, GRPO optimizes $\mathcal{J}_{\text{GRPO}}(\theta)=\mathbb{E}_{x\sim\mathcal{D}}\left[\mathcal{J}_{\text{GRPO}}(\theta,x)\right],$ where:
@@
Numerous variants have been proposed to improve GRPO, and any RLVR algorithm can be applied to our proposed TFPI stage. In all our experiments, we adopt one widely used variant, DAPO (yu2025dapo), as our RLVR algorithm for fair comparison. Details are provided in Appendix A.1.
@@ 3.1 Thinking-Free Mode Enables More Efficient Reasoning
To generate a response $y\sim\pi_{\theta}(\cdot\mid x)$ for a query $x$ using an SFT-distilled LRM $\pi_{\theta}$, the query is typically formatted with a chat template. Template 1 illustrates the template adopted by the Qwen model family (yang2025qwen3). We define ThinkingFree as an operator that transforms an input query $x$ into a modified query $x^{\prime}=\textit{ThinkingFree}{}(x)$, in which the thinking content is explicitly omitted. Under this transformation, response generation follows $y\sim\pi_{\theta}(\cdot\mid x^{\prime})$. This mechanism provides explicit control over whether reasoning content is present or absent in the generated output (see Template 2). Additional examples using other chat templates are provided in Appendix A.2. Hereinafter, we use $x$ to denote a query formatted with the thinking template (e.g., Template 1), and $x^{\prime}$ or $\textit{ThinkingFree}(x)$ to denote the corresponding thinking-free version (e.g., Template 2).
@@
Figure 2: Results of the meta-experiment on the ThinkingFree operation. Left: Average output tokens in thinking mode and ThinkingFree mode on AIME25. Right: Evolution of avg@32 and average output tokens on AIME24 with thinking-mode evaluation over training steps under 4K training response length.
@@ 3.2 Thinking-Free Training Is Beneficial to Slow-Thinking
To train an SFT-distilled LRM, the training response length should not be too short (setlur2025e3), as this is detrimental to testing performance (Polaris2025). As evidenced by the dotted line in Figure 2 (Right), training Qwen-3-4B with a 4K response length using DAPO indeed leads to a substantial drop in performance on AIME25. Given that using the ThinkingFree variant for inference can significantly reduce token consumption, we pose an audacious question: can we apply the ThinkingFree operation to all input queries during the rollout stage of RLVR? Moreover, will this approach be beneficial to preserving the original slow-thinking capability of the trained LLM?
@@ 3.3 Thinking Free Policy Initialization
hu2025prepretaining introduce a pre-pretraining stage using formal language to accelerate convergence of pretraining, while wang2025octothinker propose a mid-training stage between pretraining and RL-zero to facilitate RLVR. Inspired by these works, we ask: can a dedicated stage for SFT-distilled LRMs improve the efficiency and effectiveness of standard RLVR scaling? ThinkingFree RL, which enhances slow-thinking within short training context windows, requires substantially less compute than standard RLVR. Initializing RLVR with a ThinkingFree RL policy could therefore reduce rollout tokens while achieving stronger downstream performance. We term this step as Thinking Free Policy Initialization (TFPI), a stage preceding standard RLVR for SFT-distilled LRMs that aims to lower rollout costs, raise the ceiling of reasoning ability, and accelerate convergence.
@@
where $r(x^{\prime},y)=r(x,y)$ since the ThinkingFree operator does not alter the ground-truth answer of the original problem. In our experiments, we instantiate RLVR with DAPO, i.e., $\mathcal{J}_{\text{RLVR}}(\theta)=\mathcal{J}_{\text{DAPO}}(\theta)$ (see eq. 5). The details of DAPO (yu2025dapo) algorithm is given in Appendix A.1
@@ 4.1 Experimental Settings
Baselines. We compare TFPI with direct RLVR training from an SFT-distilled LRM (“Direct RL”) under matched total training compute (Table 1). To assess TFPI as a pre-RLVR stage, we run “TFPI + RL” with similar compute and compare against “Direct RL” (Table 2), also including competitive LRMs of the same size, such as Polaris (Polaris2025), DeepScaleR (deepscaler2025), Skywork-OR1 (he2025skyworkreasoner), and so on. For efficiency analysis, we compare with several RL-based efficient reasoning baselines (Table 3) under the same evaluation for fair comparison, including TLMRE (arora2025TLMRE), AdaptThink (zhang2025adaptthink), AutoThink (tu2025AutoThink), Laser (liu2025Laser), L1Max (aggarwal2025L1), and ThinkLess (fang2025thinkless).
Evaluation Details. Our evaluation benchmarks cover: ❶ Math Reasoning: AIME24/25 and BeyondAIME (bytedance_seed_2025_beyondaime). ❷ Multi-Task Reasoning: GPQA-Diamond (rein2024gpqa). ❸ Code Generation: LiveCodeBench (jain2024livecodebench). ❹ Instruction Following: IFEval (zhou2023ifeval). Following guo2025deepseekr1, we generate multiple outputs (ranging from 4 to 32 depending on the size of the test set) per problem and report pass@1 accuracy. Note that ❶ is in-domain evaluation, and ❷ ❸ ❹ are out-of-domain evaluation. For IFEval, we report the strict prompt accuracy. All evaluation scripts are adapted from the DeepscaleR codebase (deepscaler2025), with detailed decoding parameters provided in Appendix C.3.
@@ 5.1 Why TFPI Leads to Better Thinking-Mode Inference?
Behavioural level: The learned verification behaviours after </think> during TFPI can generalize to the slow-thinking verification occurring within the <think> and </think>. The blue lines in Figure 3 show the ratio of verification steps (i.e., the number of verification steps divided by the total number of steps; see Appendix D for details) for the training set (thinking-free mode) and AIME 25 dataset (thinking mode). We observe that the verification ratio exhibits a similar trend: a rapid drop in Stage 1, followed by steady growth in Stage 2, and a sharp increase in Stage 3. Notably, the sharp decline in Stage 1 resembles an information compression process. In Stages 2 and 3, the model begins to explore more extensively (Figure 3 Right), which may explain why TFPI achieves superior performance. As verification is believed to be vital for slow-thinking reasoning (setlur2025e3), the observed generalization of verification behavior suggests a transfer from thinking-free training to inference in thinking mode.
@@ 6 Related Work
LRMs & Efficiency RLVR. RLVR has enabled the development of numerous high-performing large reasoning models (LRMs) (yang2025qwen3; zeng2025glm4_5; Polaris2025), inspiring research on model behaviors (liu2025drgrpo; wang2025beyond8020), novel algorithms (liu2025drgrpo; zheng2025gspo), multimodal extensions (meng2025mmeureka; xiao2025perception; wang2025papo), tool integration (jin2025searchr1; feng2025retool; li2025torl; xue2025simpletir; team2025kimik2), and other related directions (xu2025doublechecker; zhang2025verifyrl; chen2025verithinker; zhang2025critiquegrpo). RLVR can be applied directly to base LLMs (“RL zero”) (zeng2025simplerl; guo2025deepseekr1) or initialized from SFT-distilled long-CoT models, the latter typically yielding stronger results (deepscaler2025; Polaris2025). A major challenge for RLVR is the cost of training with long contexts, as longer outputs are often necessary for harder tasks (shrivastava2025gfpo; zeng2025glm4_5), consuming high computational resources. Multi-stage RLVR mitigates this by starting with shorter contexts and gradually extending them (deepscaler2025; Polaris2025; he2025skyworkreasoner), while algorithmic approaches modify GRPO to reduce length bias (yu2025dapo; liu2025drgrpo; wang2025beyond8020). Orthogonal to these strategies, we introduce TFPI as a lightweight stage before RLVR, improving efficiency and strengthening the slow-thinking mode at inference with minimal training cost, thus facilitating more effective subsequent RLVR.
Efficient Reasoning. To address the issue of overthinking (chen2024overthinking; sui2025overthinkingsurvey), considerable efforts have been made, including prompt-based methods (muennighoff2025s1; yang2025deer; fu2025dynasor; chen2025seal; fu2025deepconf), SFT-based approaches (kang2025c3ot; ma2025cot-value; munkhbat2025self-train-efficient; luo2025o1-pruner), and RL-related designs. RL-related approaches can be further categorized into length-based reward shaping (team2025kimik1.5; aggarwal2025L1; arora2025TLMRE; liu2025Laser), integration of fast and slow thinking (fang2025thinkless; zhang2025adaptthink; lou2025adacot; tu2025AutoThink; jiang2025LHRMs; zhang2025ASRR), and thinking budget control (li2025selfbudgeter; hammoud2025curriculum-grpo; wen2025budgetthinker). These methods primarily trade accuracy for efficiency and rely on specialized reward functions or training strategies to encourage more efficient reasoning. In contrast, our proposed TFPI naturally yields even more efficient LRMs without specialized rewards or training designs.
@@ A.1 RLVR Algorithms
Numerous variants have been proposed to improve GRPO. For example, DAPO (yu2025dapo) introduces token-level normalization and dynamic sampling; Dr GRPO (liu2025drgrpo) removes length bias to prevent incorrect responses from growing longer over time; and wang2025beyond8020 train selectively on forking tokens (see also Section 6). Our TFPI is orthogonal to these RLVR algorithms. That is to say, any RLVR algorithm can be applied to our proposed TFPI stage. To mitigate the effect of RLVR algorithms, we employ DAPO as our RLVR algorithm in all experiments for fair comparison.
@@
Building on GRPO, DAPO (yu2025dapo) introduces a clip-higher mechanism, incorporates dynamic sampling and applies a token-level policy gradient loss. The objective is given by $\mathcal{J}_{\text{DAPO}}(\theta)=\mathbb{E}_{x\sim\mathcal{D}}\left[\mathcal{J}_{\text{DAPO}}(\theta,x)\right],$ where:
@@ A.2 ThinkingFree Operation
In Section 3.1, we have shown chat templates for both the original query $x$ and its thinking-free version $x^{\prime}=\textit{ThinkingFree}(x)$ for Qwen models. Here, we showcase one additional example under the DeepSeek (guo2025deepseekr1) template as below.
@@ B.1 Token Consumption of ThinkingFree
For DS-1.5B, we use the decoding parameters suggested by guo2025deepseekr1 in thinking mode: temperature = 0.6, top-$p$ = 0.95, and top-$k$ = 1-1. For Qwen3-4B, we adopt the recommended settings from yang2025qwen3 in thinking mode: temperature = 0.6, top-$p$ = 0.95, and top-$k$ = $20$. For ThinkingFree, we set temperature = 0.7, top-$p$ = 0.8, and use both top-$k$ = 1-1 and top-$k$ = $20$. The maximum output length is fixed at 32K tokens for both modes. For evaluation, we sample 32 generations per query on AIME 2025 and report the average number of output tokens. The parameters are given in Table 4 and the results are shown in Figure 2 (Left).